OpenStudy (anonymous):

A company produces baseballs at 3 different plants: plant A, plant B and plant C. doubling plant A's production is equal to 1/3 of the company's total production. doubling plant B's production is equal to the company's total production.tripling plant C's production is equal to the company's total production what us the ratio of plant A's production to plant B's production to plant C's production?

OpenStudy (anonymous):

2*A=(1/3)Total 2*B=Total 3*C=Total solving for T and setting equal 6A=2B=3C

OpenStudy (anonymous):

sorry i couldn't get u...will you please elaborate it...